Activity during this period has been low with probable continued growth of the new dome within the scar left by the December 26 collapse. Seismic activity has been relatively low and dominated by rockfalls.
The predominant activity throughout this period has been small rockfalls and pyroclastic flows from the new dome growing in the scar formed by the December 26 collapse. The rockfalls mostly occurred in the upper part of the White River, but some small rockfalls were also seen down the Tar River valley; talus material has been accumulating in the depression outwith the remains of Galway's Wall. The rockfalls produced dilute ash clouds which generally drifted westwards. A brief helicopter flight around the south-west of the volcano on 12 January revealed glimpses into the collapse scar from the Boxing Day event, where a new dome with a steep front face and extensive talus slope had grown. Clear weather on 12 and 13 January showed that the volcano was continuously venting ash and steam, with the top of the vertical column reaching 8 to 10,000 feet. Activity on other sectors of the dome, apart from the December 26 growth, has been low.
Heavy rain on 11 January resulted in hot mudflows being generated in the Belham River valley. Large logs up to 6 m long and rocks up to 60 cm diameter were transported with the muddy water which was about 10 m wide at the Belham bridge. Temperatures of the water about one hour after the peak in activity were only 27.4 degrees centigrade (approximately 9 degrees above air temperature), but rocks out of the main flow on the golf course were still steaming and were hot to touch. Erosion of the recent deposits by the swollen river near Molyneaux produced small ash clouds as the sides of the river valley collapsed into the water.
There is an ongoing investigation into the deposits from the December 26 event. The observations, measurements and preliminary interpretations will be reported in an MVO Special Scientific Report (Special Report 6: "The Boxing Day Collapse, 26 December 1997"), and are summarised here.
Observations confirmed that the areas of Trials and Fairfield had both been badly affected by the 26 December event. In Trials, most buildings were damaged. Roofs have collapsed and there is some fire damage but most structures are still standing. In Fairfield, the roofs of some houses have collapsed as a result of heavy ash fall, but the pyroclastic surge probably did not reach this far, as there is little fire damage. The villages of St. Patrick's and Morris' have been almost completely destroyed with only a few foundations of buildings remaining.
Several visits to localities in the devastated area were possible. Samples were taken, sections were dug and temperatures of the deposits have been measured (MVO Special Report 6: "The Boxing Day Collapse, 26 December 1997"). There are three main types of deposit: debris avalanche deposits, pyroclastic flow and ash cloud deposits (including co-ignimbrite ash), and blast deposits. There is also considerable erosion of some surfaces. The relationships between different facies are not simple, but it is thought that a sector collapse occurred. This involved slippage of material from the area around Galway's Soufriere (which was probably weakened by hydrothermal activity), part of Galway's Wall and new talus and dome rock. This then triggered a large conventional dome collapse with associated pyroclastic flows and ash cloud surges, and culminated in an energetic lateral blast. Work is still ongoing to test these preliminary conclusions.
For the whole of this period seismicity was relatively low and dominated by rockfall signals. There were a number of isolated hybrid swarms and one small volcano-tectonic earthquake swarm but these were not followed by any observed increase in surface activity. Throughout the reporting period there were intervals of increased seismic amplitude with the time between such intervals varying from 6 to 24 hours. This increased amplitude was recorded on all stations even when no events were being recorded - although most rockfall signals were recorded during these intervals and contributed to the amplitude increase. The frequency content of the background seismicity was the same at these times as at times when the amplitude was lower, making it difficult to interpret this feature of the data.
Table 1: Earthquake types
These earthquake counts are of events that triggered the broadband network's event recording system between 00:00 and 00:00 each day (local time).
Date VT Hybrid LP Dome RF LPRF* HYRF* 04 Jan 98 0 8 3 68 1 2 05 Jan 98 1 25 12 88 6 0 06 Jan 98 1 50 33 22 3 1 07 Jan 98 1 40 9 66 2 1 08 Jan 98 7 23 13 29 0 0 09 Jan 98 10 17 1 29 0 0 10 Jan 98 14 16 3 33 1 0 11Jan 98 1 14 6 39 4 2 12 Jan 98 8 15 5 28 0 1 13 Jan 98 1 14 3 28 0 0 14 Jan 98 4 5 3 17 1 0 15 Jan 98 4 19 1 28 0 0 16 Jan 98 5 17 10 15 1 1 17 Jan 98 3 13 4 19 0 0
* LPRF: LP earthquake followed by rockfall signal. HYRF: Hybrid earthquake followed by rockfall signal. The LPs, hybrids and rockfalls in these signals are also counted in their respective columns.
Table 2: Swarms
Start Duration Hybrid LP VT 7 Jan 98 02:23 1.73 15 3 0 7 Jan 98 23:08 0.88 10 0 0 8 Jan 98 15:17 2.68 15 1 0 10 Jan 98 13:06 2.57 3 0 8 16 Jan 98 03:25 2.05 11 0 0
The GPS networks BIGNET (Harris, Whites, Long Ground, Windy Hill and Broderick's) and LEESNET (Old Towne, Waterworks, St Georges Hill and Lees Yard) were occupied during the reporting period. No clear trends have been discerned in the data for LEESNET yet. Some of the stations in BIGNET continue to show slow movements. Long Ground and Whites show continued movement to the NE and N respectively. The Harris-Windy Hill baseline appears to have undergone two lengthening-shortening cycles since the measurements began in June 1996. The first cycle ended about the middle of May 1997 involving a lengthening and shortening of 2 cm. The second cycle involved lengthening and shortening by almost 4 cm and the line is close to its May 1997 length. Long occupations of the stations at Hermitage and Tar River were made running a base station at Harris. Hermitage shows continued movement to the NNE at around 0.3 cm per week. Tar River had not been occupied since 6 March 1997. Since that time it has moved 5 cm to the NNE.
A volume survey of the deposits in the upper reaches of the White River Valley was carried out on 17 January.
The data from this survey were combined with that of the previous survey, on 4 January, of the lower part of the valley and the delta. This gave a total volume for the deposits in the White River valley of 45.9 million m3. Included in this is an estimate for the volume of the surge component, covering an area of 9.1 km2 , which was calculated to be 1.8-3.2 million m3.
Scar volumes have been estimated using data generated from construction of cross-sections and assuming relatively simple geometries. A breakdown of this reveals that material lost during the collapse of 26 Dec comprises 20 million m3 of hydrothermally altered Galway's Soufriere rock, 5 million m3 of Galways wall material, 26 million m3 of the November lava dome and 26 million m3 of dome talus.
From the figures above it is calculated that 20 million m3 has been lost to sea which is consistent with the size of the tsunami wave that was generated as a result.
The volume of this event at 46 million m3 is 5 times the magnitude of largest dome collapse event of the eruption to date. The largest single event prior to this was on 21 September 1997 with 9 million m3.
Dust Trak monitoring has been carried out at four fixed sites. Each value is an average PM10 concentration over approximately 24 hours* and have been low over this period.
*13th January - Dust Traks were taken in to be downloaded, value is an average for approximately a 15 hour period. Table 3 : PM10 (concentration in mg/m3)
These values are approximate 24 hr averages of the PM10 concentration.
Site 04-Jan 05-Jan 06-Jan 07-Jan 08-Jan 09-Jan 10-Jan Mango 0.010 0.011 0.013 0.008 0.007 0.006 0.015 CPS 0.017 0.024 0.039 0.024 0.023 0.021 0.016 St. Peter's 0.013 0.024 0.014 0.010 0.011 0.012 0.015 MVO north 0.000 0.000 0.000 0.000 0.000 0.000 0.000 Site 11-Jan 12-Jan 13-Jan 14-Jan 15-Jan 16-Jan 17-Jan Mango 0.010 0.016 0.015 0.006 0.009 0.007 0.009 CPS 0.011 0.049 0.015 0.038 0.044 0.042 0.036 St. Peter's 0.012 0.020 0.015 0.011 0.006 0.009 0.012 MVO north 0.000 0.000 0.001 0.002 - - - 24 Hour averages: <0.05 mg m-3 Low 0.05-0.1 mg m-3 Raised 0.1-0.3 mg m-3 Very High >0.3 mg m-3 Alert Locations: St. Peter's: The Dust Trak is outside on the balcony of a villa in St Peter's, called St. Peter's Place. CPS: Catholic Primary School The Dust Trak is outside in the area under the roof where some of the children have classes. The school is in Palm Loop near Woodlands. MVO north: The new Volcano observatory on Mongo Hill near St.John's, the Dust trak is on the third story of the building with the sampling tube stuck out of the window. Mango: The Dust Trak is outside on the verandah of a villa in Mango Drive in Woodlands
MVO Staff Changes
Mr. Richie Robertson (Seismic Research Unit)
Dr. Simon Young (British Geological Survey)
Prof. Steve Sparks (Bristol University)
Dr. Jenni Barclay (University of California, Berkeley)
Mr. Thomas Christopher (MVO)
Ms. Caroline Choux (Blaise-Pascal University, Clermont Ferrand)
Prof. Barry Voight (Penn State University)
Mr. Graham Ryan (Lancaster University)
Ms. Chloe Harford (Bristol University)
Ms. Eliza Calder (Bristol University)
As of 18 January, Richie Robertson (SRU) is the current chief scientist of MVO. Gill Norton (BGS) is the deputy chief scientist.
Other overseas MVO staff are:
Dr. Simon Young (BGS)
Dr. Richard Herd (BGS)
Dr. Richard Luckett (BGS)
Mr. Lutchman Pollard (SRU)
Ms. Lucy Ritchie (Luton University)
Ms. Hayley Duffell (BGS)